# Learning convolution operators on compact abelian groups

### v2

@@ Abstract.
We consider the problem of learning convolution operators associated to compact Abelian groups. We study a regularization-based approach and provide corresponding learning guarantees, discussing natural regularity condition on the convolution kernel. More precisely, we assume the convolution kernel is a function in a translation invariant Hilbert space and analyze a natural ridge regression (RR) estimator. Building on existing results for RR, we characterize the accuracy of the estimator in terms of finite sample bounds. Interestingly, regularity assumptions which are classical in the analysis of RR, have a novel and natural interpretation in terms of space/frequency localization. Theoretical results are illustrated by numerical simulations.
@@ 1. Introduction
The key problem in machine learning is estimating an input/output function $f$ of interest from random input/output pairs $(x_{i},y_{i})_{i=1}^{n}$. Classically, inputs are vectors in $\mathbb{R}^{d}$, and outputs are binary or scalar values. However, as the scope and number of machine learning applications expand frantically, it is interesting to consider functional relationships between more general inputs and outputs. Relevant to this study is the case where both inputs and outputs are elements of infinite-dimensional spaces, such as Hilbert or Banach spaces, so that $f$ can be viewed as an operator. This setting has recently received considerable attention, driven by applications in image and signal processing, and more generally in scientific and engineering contexts where data are described by integral and partial differential equations (PDEs); see e.g.[23] and references therein.
In this paper, we focus on a special class of linear operators, namely convolution operators on an Abelian group $G$:
@@ 1.1. Related work
Operator learning. As already mentioned, operator learning has received a lot of attention lately. On the one hand, there has been a growing literature on neural network approaches, the so-called neural operators [22]; see [16] and the references therein for an overview. On the other hand, learning-theoretic studies have largely focused on learning linear operators with kernel methods, since they are amenable to a more complete theoretical analysis. A recent survey can be found in [23], whereas a partial list of references includes [11, 13, 43, 30, 21, 1, 3, 42]. These latter studies consider estimators and technical tools analogous to those in this paper. Indeed, the convolution operators we consider are an example of linear operators. However, their special structure allows for a tailored analysis that highlights the specific structure of the problem. In particular, it is possible to focus on the convolution kernel as the primary object of interest, using functional analytic tools rather than operator analytic tools, and learning bounds in different norms can be derived, leveraging the regularity properties of the convolution kernel. In addition, regularity assumptions needed to derive learning bounds can be related to localization properties.
@@
Learning Green functions. Motivated by applications in PDE, another class of linear operators has been considered, namely integral kernel operators where the kernel can be related to Green functions; see, e.g., [4] and references therein. In our setting, the hypothesis space has a special structure, since it is defined by a translational invariant kernel and reflects the group structure. The analysis in this case cannot be recovered from the case of general integral kernel operators.
@@
Functional regression. The problem we study is related to functional data analysis, where input/output data are represented as functions in a continuous domain; see, e.g., [33, 28] and [10, 19, 34]. For example, the studies in [26, 25, 46] address the analysis within Reproducing Kernel Hilbert spaces. The framework we consider is more general in view of the general Abelian group setting. It is more specific since we focus on convolution kernels, and we can derive and interpret results in different norms.
@@
Linear time-invariant system estimation. Learning convolution kernels and convolution operators is close to the central identification of linear time-invariants in control theory and signal processing. We refer to [27] for a classic reference and to [5] for modern data-driven approaches that use optimization and machine learning. Estimating linear time-invariant systems involves learning the so-called impulse response or transfer function, which can often be represented as a convolution operator. Indeed, the tools we employ, such commutative Harmonic Analysis, parallel the techniques used in LTI system analysis. Our contributions extend these ideas by focusing on a more general framework where the underlying structure is defined by a compact Abelian group, allowing for novel insights and guarantees in a broader stochastic setting. Moreover, we consider random inputs and additive noise in the observations, making the problem inherently statistical.
@@
Blind deconvolution. The problem of learning a convolution kernel is related to the so-called blind deconvolution; see, e.g., [20]. A number of machine learning approaches primarily based on neural networks have been considered for this problem [41, 40, 14, 12, 6, 24], particularly with respect to the problem of image recovery from blurred photographs [9, 45]. Typically, only a discrete setting is considered, and no theoretical results are developed. Closely related to our study is the approach in [41], which focused on algorithmic aspects. Here, we complement these results by considering a continuous setting and deriving learning theory guarantees.
@@ 1.2. Contribution
In the context of the learning model in (1), our main contribution is the analysis of the learning properties of a ridge regression estimator in terms of non-asymptotic learning bounds. More precisely, we consider convolution operators defined by a compact Abelian group $G$. With this choice, we identify the input and output spaces with $L^{1}(G)$ and $L^{2}(G)$, respectively, which are the Lebesgue spaces defined by the Haar measure associated with $G$. We further assume that the convolution kernel $w^{*}$ in (1) belongs to $L^{2}(G)$, so that the convolution operator is well defined from $L^{1}(G)$ into $L^{2}(G)$.
@@
The remainder of the paper is organized as follows. In Section 2, we provide essential background on group theory and Harmonic Analysis. Section 3 introduces our statistical model and methodology for learning convolution operators, including the use of translation-invariant Hilbert spaces and ridge regression. In Section 4, we present the theoretical analysis of the learning error, discuss the main results, and explore the implications of a-priori assumptions related to space and frequency localization. Section 5 is devoted to numerical simulations, where we validate the theoretical error bounds and demonstrate the applicability of our framework to approximating heat kernels in the context of partial differential equations. We conclude the paper in Section 6 with a summary of our contributions and directions for future research. Detailed proofs of our results are provided in the appendices.
@@ 2. Background on group theory and Harmonic Analysis
In this section, we recall basic notions and facts from group theory and from commutative Harmonic Analysis. In particular, we focus on compact Abelian groups and introduce the notions of convolution operators and Fourier transform, which we illustrate with some examples. We refer to some standard reference such as [38] for further reading.
@@
$(\mathcal{F}^{-1}\widehat{x})(t)=\sum_{\xi\in\widehat{G}}\widehat{x}_{\xi}\,\langle{\xi},{t}\rangle\qquad t\in G,$
@@
$\lVert{x\convolution y}\rVert_{p}\leq\lVert{x}\rVert_{1}\lVert{y}\rVert_{p}\ ,$ (4)
@@
is jointly continuous.
Moreover, if $x\in L^{1}$, $y\in L^{p}$ and $z\in L^{q}$ with $11/p+1/q=$ 1,
@@
$e_{\xi}(t)=\langle{\xi},{t}\rangle\qquad t\in G\qquad\qquad(\widehat{e}_{\xi})_{\xi^{\prime}}=\delta_{\xi,\xi^{\prime}}\qquad\xi^{\prime}\in\widehat{G},$
@@
$\begin{split}&\mathcal{F}e_{\xi}=\widehat{e}_{\xi}\qquad\qquad\xi\in\widehat{G}\\ &x\convolution e_{\xi}=(\mathcal{F}x)_{\xi}\,e_{\xi}\qquad\xi\in\widehat{G}\\ &\langle{y},{e_{\xi}}\rangle_{2}=(\mathcal{F}y)_{\xi}\qquad y\in L^{2},\xi\in\widehat{G}\end{split}\quad.$ (8)
@@
The indexing ${x}=({x}_{0},\dots,{x}_{n-1})$ is possible because the indices are evaluated mod $n$. As vector spaces, $L^{P}=\ell^{p}=\mathbb{C}^{N}$, the convolution operator
$C_{y}:\mathbb{C}^{N}\to\mathbb{C}^{N}\qquad C_{y}x=x\convolution y$
is the $N\times N$ circulant matrix
@@ 3.1. Statistical model
We let $\mathcal{X}=L^{1}$ and $\mathcal{Y}=L^{2}$ and we view $L^{1}$ as the input space and $L^{2}$ as the output space. We let $\mathcal{H}\subseteq L^{2}$ be a Hilbert space of hypothesis and $j:\mathcal{H}\to L^{2}$ the canonical embedding. Let $(X,Y)$ be a pair of random variables such that
@@
    where $C_{*}$ is the convolution operator
    $C_{*}:L^{1}\to L^{2}\qquad C_{*}x=x\convolution{j(w_{*})}\quad\text{for some }w_{*}\in\mathcal{H},$ (13)
    and $\epsilon$ is a random variable in $L^{2}$ such that
@@
Eqs. (12)–(14) describe a natural regression model: the random output $Y$ is a noisy image of a convolution operator of the random input $X$ for an unknown convolution kernel $w_{*}$ belonging to the hypothesis space $\mathcal{H}$. This last assumption in the context of a regression problem corresponds to the assumption that the model is well-specified.
@@ 3.2. Translation invariant Hilbert spaces
We now describe the class of hypothesis spaces under consideration. Given the properties of convolution operators, a natural choice is Hilbert spaces invariant under translations by any elements $t\in G$. These spaces can be easily characterized using the Fourier transform, as we recall next. We refer to [2], for further details. Indeed, let $\{\widehat{K}\}_{\xi\in\widehat{G}}$ be a family such that
$0\leq\widehat{K}_{\xi}\leq D^{2}_{K}\qquad\xi\in\widehat{G}$ (18)
@@
It is a standard result that $\mathcal{H}$ is a Hilbert space with a continuous embedding $j:\mathcal{H}\to L^{2}$. Moreover, $\mathcal{H}$ is invariant under translations, i.e $T_{t}\mathcal{H}=\mathcal{H}$ for any $t\in G$.
@@
$\begin{split}&\widehat{G}_{*}=\{\xi\in\widehat{G},\widehat{K}_{\xi}\neq 0\}\\ &f_{\xi}=\widehat{K}_{\xi}^{\frac{1}{2}}e_{\xi}\qquad\xi\in\widehat{G}_{*}\end{split}\ ,$ (21)
@@
We recall that, given an arbitrary set $G$, a reproducing kernel Hilbert space $\mathcal{H}$ is a Hilbert space of functions from $G$ to $\mathbb{R}$, equipped with a kernel function $k:G\times G\to\mathbb{C}$ such that, for all $t\in G$, $k(t,\cdot)\in\mathcal{H}$, and for all $w\in\mathcal{H}$, $w(t)=\langle{w},{k(t,\cdot)}\rangle_{\mathcal{H}}$. From this definition, it follows that $k$ is positive definite. It is a classic fact that the converse also holds: every positive definite kernel uniquely defines a reproducing kernel Hilbert space.
@@
where $\zeta$ is the Riemann zeta function, then $\mathcal{H}$ is the Sobolev space $H^{s}$ of periodic functions on $\mathbb{R}$ with period $1$, and $\mathcal{H}$ is a dense subspace of $L^{2}$. If $s>1/2$, $\mathcal{H}$ is a reproducing kernel Hilbert space. With the choice $s=1$, i.e
@@ 3.3. Ridge regression
We next discuss a natural kernel ridge regression estimator adapted to learn convolution operators. As shown next some specific expressions are available in this case.
@@
$(\mathcal{F}j(w_{n}^{\lambda}))_{\xi}=\begin{cases}\dfrac{\frac{1}{n}\sum_{i=1}^{n}(\mathcal{F}Y_{i})_{\xi}\overline{(\mathcal{F}X_{i})_{\xi}}}{\frac{1}{n}\sum_{i=1}^{n}|(\mathcal{F}X_{i})_{\xi}|^{2}+\lambda\widehat{K}_{\xi}^{-1}}&\xi\in\widehat{G}_{*}\\ 0&\xi\notin\widehat{G}_{*}\end{cases}\,.$ (32)
@@ 4. Theoretical analysis of the learning error
where the second equality is due to the (12), (14) and
@@
where $\Sigma:\mathcal{H}\to\mathcal{H}$ is the diagonal operator on the basis $\{f_{\xi}\}_{\xi\in\widehat{G}_{*}}$ defined by (21), i.e.
@@
Assume that $G=\mathbb{R}/\mathbb{Z}$. In Example 3 we have that
@@
is yet another error measure which is more stringent than either (33) or (34).

### v3

@@ Abstract.
We consider the problem of learning convolution operators associated to compact Abelian groups. We study a regularization-based approach and provide corresponding learning guarantees under natural regularity conditions on the convolution kernel. More precisely, we assume the convolution kernel is a function in a translation invariant Hilbert space and analyze a natural ridge regression (RR) estimator. Building on existing results for RR, we characterize the accuracy of the estimator in terms of finite sample bounds. Interestingly, regularity assumptions which are classical in the analysis of RR, have a novel and natural interpretation in terms of space/frequency localization. Theoretical results are illustrated by numerical simulations.
@@ 1. Introduction
The key problem in machine learning is estimating an input/output function $f$ of interest from random input/output pairs $(x_{i},y_{i})_{i=1}^{n}$. Classically, inputs are vectors in $\mathbb{R}^{d}$, and outputs are binary or scalar values. However, as the scope and number of machine learning applications expand frantically, it is interesting to consider functional relationships between more general inputs and outputs. Relevant to this study is the case where both inputs and outputs are elements of infinite-dimensional spaces, such as Hilbert or Banach spaces, so that $f$ can be viewed as an operator. This setting has recently received considerable attention, driven by applications in image and signal processing, and more generally in scientific and engineering contexts where data are described by integral and partial differential equations (PDEs), see [23] and references therein.
In this paper, we focus on a special class of linear operators, namely convolution operators on an Abelian group $G$,
@@ 1.1. Related work
Operator learning. As already mentioned, operator learning has received significant attention lately. On the one hand, there has been a growing literature on neural network approaches, such as the so-called neural operators [22]; see also [16] and the references therein for an overview. On the other hand, learning-theoretic studies have largely focused on learning linear operators with kernel methods, since they are amenable to a more complete analysis. A recent survey can be found in [23], whereas a partial list of references includes [11, 13, 44, 30, 21, 1, 3, 43]. These latter studies consider estimators and technical tools analogous to those in this paper. Indeed, the convolution operators we consider are an example of linear operators. However, their special structure allows for a tailored analysis that highlights the specific structure of the problem. In particular, it is possible to focus on the convolution kernel as the primary object of interest, using functional analytic tools rather than operator analytic tools, and learning bounds in different norms can be derived, leveraging the regularity properties of the convolution kernel. A recent contribution in this direction is [48], which considers a more general class of integral operators and derives minimax rates in certain spectral Sobolev spaces. While their approach covers many integral kernels, our setting leverages the commutative group structure and frequency-domain localization, leading to interpretations that are not considered in their framework. In particular, regularity assumptions needed to derive learning bounds in our setting can be related to localization properties.
@@
Learning Green functions. Motivated by applications in PDEs, another class of linear operators that has been considered are integral kernel operators, where the kernel can be related to Green functions; see [4] and references therein. In our setting, the hypothesis space has a special structure, since it is defined by a translational invariant kernel and reflects the group structure. The analysis in this case cannot be recovered from the case of general integral kernel operators.
@@
Functional regression. The problem we study is related to functional data analysis, where input/output data are represented as functions in a continuous domain; see, e.g., [33, 28] and [10, 19, 34]. For example, the studies in [26, 25, 47] address the analysis within Reproducing Kernel Hilbert spaces. Our framework is more general because we consider the general Abelian group setting. At the same time, it is more specific since we focus on convolution kernels, and we can derive and interpret results in different norms.
@@
Linear time-invariant system estimation. Learning convolution kernels and convolution operators is close to identification of linear time-invariants in control theory and signal processing. We refer to [27] for a classic reference and to [5] for modern data-driven approaches that use optimization and machine learning. Estimating linear time-invariant systems involves learning the so-called impulse response or transfer function, which can often be represented as a convolution operator. Indeed, the tools we employ, such as commutative Harmonic Analysis, parallel the techniques used in identification of linear time-invariant systems. Our contributions extend these ideas by focusing on a more general framework where the underlying structure is defined by a compact Abelian group. Moreover, we consider random inputs and additive noise in the observations, making the problem inherently statistical.
@@
Blind deconvolution. The problem of learning a convolution kernel is related to the so-called blind deconvolution problem; see, e.g., [20]. A number of machine learning approaches primarily based on neural networks have been considered for this problem [41, 40, 14, 12, 6, 24], particularly with respect to the problem of image recovery from blurred photographs [9, 46]. Typically, only a discrete setting is considered, and no theoretical results are developed. Closely related to our study is the approach in [41], which focused on algorithmic aspects. Here, we complement these latter results by considering a continuous setting and deriving learning theoretic guarantees.
@@ 1.2. Contribution
In the context of the learning model in eq. 1, our main contribution is the analysis of the learning properties of a ridge regression estimator in terms of non-asymptotic learning bounds. More precisely, we consider convolution operators defined by a compact Abelian group $G$. With this choice, we identify the input and output spaces with $L^{1}(G)$ and $L^{2}(G)$, respectively, which are the Lebesgue spaces defined by the Haar measure associated with $G$. We further assume that the convolution kernel $w^{*}$ in eq. 1 belongs to $L^{2}(G)$, so that the convolution operator is well defined from $L^{1}(G)$ into $L^{2}(G)$.
@@
The remainder of the paper is organized as follows. In Section 2, we provide essential background on group theory and Harmonic Analysis. Section 3 introduces our statistical model and methodology for learning convolution operators, including the use of translation-invariant Hilbert spaces and ridge regression. In Section 4, we present the theoretical analysis of the learning error, discuss the main results, and explore the implications of a-priori assumptions in terms of space and frequency localization properties. Section 5 is devoted to numerical simulations, where we validate the theoretical error bounds and demonstrate the applicability of our framework to approximating heat kernels in the context of partial differential equations. We conclude the paper in Section 6 with a summary of our contributions and directions for future research. Detailed proofs of our results are provided in the appendices.
@@ 2. Background on group theory and Harmonic Analysis
In this section, we recall basic notions and facts from group theory and from commutative Harmonic Analysis. In particular, we focus on compact Abelian groups and introduce the notions of convolution operators and Fourier transform, which we illustrate with some examples. We refer to some standard reference such as [38] for further readings.
@@
$(\mathcal{F}^{-1}\widehat{x})(t)=\sum_{\xi\in\widehat{G}}\widehat{x}_{\xi}\,\langle{\xi},{t}\rangle,\qquad t\in G,$
@@
$\lVert{x\convolution y}\rVert_{p}\leq\lVert{x}\rVert_{1}\lVert{y}\rVert_{p}$ (4)
@@
is jointly continuous. For any fixed $y\in L^{p}$, we define the convolution operator
$C_{y}:L^{1}\to L^{p},\quad(C_{y}x)(t)=x\convolution y(t),\quad t\in G.$
The inequality (4) implies that $C_{y}$ is a bounded linear operator.
If $x\in L^{1}$, $y\in L^{p}$ and $z\in L^{q}$ with $11/p+1/q=$ 1,
@@
$e_{\xi}(t)=\langle{\xi},{t}\rangle,\qquad t\in G,\qquad\qquad(\widehat{e}_{\xi})_{\xi^{\prime}}=\delta_{\xi,\xi^{\prime}},\qquad\xi^{\prime}\in\widehat{G},$
@@
$\begin{split}&\mathcal{F}e_{\xi}=\widehat{e}_{\xi},\qquad\qquad\xi\in\widehat{G}\\ &x\convolution e_{\xi}=(\mathcal{F}x)_{\xi}\ e_{\xi},\qquad\xi\in\widehat{G}\\ &\langle{y},{e_{\xi}}\rangle_{2}=(\mathcal{F}y)_{\xi},\qquad y\in L^{2},\ \xi\in\widehat{G}.\end{split}\quad$ (8)
@@
The indexing ${x}=({x}_{0},\dots,{x}_{n-1})$ is possible because the indices are evaluated mod $n$. As vector spaces, $L^{P}=\ell^{p}=\mathbb{C}^{N}$, the convolution operator $C_{y}:\mathbb{C}^{N}\to\mathbb{C}^{N},\ C_{y}x=x\convolution y$ is the $N\times N$ circulant matrix
@@ 3.1. Statistical model
We let $\mathcal{X}=L^{1}$ and $\mathcal{Y}=L^{2}$ and we view $L^{1}$ as the input space and $L^{2}$ as the output space. We let $\mathcal{H}\subseteq L^{2}$ be a Hilbert space of hypothesis and $j:\mathcal{H}\hookrightarrow L^{2}$ the canonical embedding. Let $(X,Y)$ be a pair of random variables such that
@@
    where $C_{*}:=C_{w_{*}}$, with $C_{\ast}x=x\convolution{j(w_{*})}$ for some $w_{*}\in\mathcal{H}$, and and $\epsilon$ is a random variable in $L^{2}$ such that
@@
Eqs. (11)–(13) describe a natural regression model: the random output $Y$ is a noisy image of the random input $X$ convolved with an unknown convolution kernel $w_{*}$ belonging to $\mathcal{H}$. This last assumption corresponds to the assumption that the model is well-specified in the context of regression.
@@ 3.2. Translation invariant Hilbert spaces
We now describe the class of hypothesis spaces we consider. Given the properties of convolution operators, a natural choice is Hilbert spaces invariant under translations by any elements $t\in G$. These spaces can be easily characterized using the Fourier transform, as we recall next. We refer to [2], for further details. Indeed, let $\{\widehat{K}\}_{\xi\in\widehat{G}}$ be a family such that
$0\leq\widehat{K}_{\xi}\leq D^{2}_{K},\qquad\xi\in\widehat{G},$ (17)
@@
It is a standard result that $\mathcal{H}$ is a Hilbert space with a continuous embedding $j:\mathcal{H}\to L^{2}$. Moreover, $\mathcal{H}$ is invariant under translations, i.e. $T_{t}\mathcal{H}=\mathcal{H}$ for any $t\in G$.
@@
$\begin{split}&\widehat{G}_{*}=\{\xi\in\widehat{G},\widehat{K}_{\xi}\neq 0\}\\ &f_{\xi}=\widehat{K}_{\xi}^{\frac{1}{2}}e_{\xi}\qquad\xi\in\widehat{G}_{*}.\end{split}$ (20)
@@
We recall that, given an arbitrary set $G$, a reproducing kernel Hilbert space $\mathcal{H}$ is a Hilbert space of functions from $G$ to $\mathbb{R}$, equipped with a kernel function $k:G\times G\to\mathbb{C}$ such that, for all $t\in G$, $k(t,\cdot)\in\mathcal{H}$, and for all $w\in\mathcal{H}$, $w(t)=\langle{w},{k(t,\cdot)}\rangle_{\mathcal{H}}$. From this definition, it follows that $k$ is positive definite. It is a classic fact that the converse also holds: every positive definite kernel uniquely defines a reproducing kernel Hilbert space [42].
@@
where $\zeta$ is the Riemann zeta function, then $\mathcal{H}$ is the Sobolev space $H^{s}$ of periodic functions on $\mathbb{R}$ with period $1$, and $\mathcal{H}$ is a dense subspace of $L^{2}$. If $s>1/2$, $\mathcal{H}$ is a reproducing kernel Hilbert space. With the choice $s=1$, i.e.
@@ 3.3. Ridge regression
We next discuss a natural ridge regression estimator adapted to learn convolution operators. As shown next, specific expressions are available in this case.
@@
$(\mathcal{F}j(w_{n}^{\lambda}))_{\xi}=\begin{cases}\dfrac{\frac{1}{n}\sum_{i=1}^{n}(\mathcal{F}Y_{i})_{\xi}\overline{(\mathcal{F}X_{i})_{\xi}}}{\frac{1}{n}\sum_{i=1}^{n}|(\mathcal{F}X_{i})_{\xi}|^{2}+\lambda\widehat{K}_{\xi}^{-1}}&\xi\in\widehat{G}_{*}\\ 0&\xi\notin\widehat{G}_{*}.\end{cases}\,$ (31)
@@ 4. Theoretical analysis of the learning error
where the second equality is due to eqs. (12) and (13), and
@@
where $\Sigma:\mathcal{H}\to\mathcal{H}$ is the diagonal operator on the basis $\{f_{\xi}\}_{\xi\in\widehat{G}_{*}}$ defined by eq. 20, i.e.
@@
Assume that $G=\mathbb{R}/\mathbb{Z}$. In Example 3, we have that
@@
is yet another error measure which is more stringent than either eq. 32 or eq. 33.